OpenStudy (anonymous):

multipy and simplify using factoring assume all expressions under radicals represent nonnegative numbers simplify answerin radical form

OpenStudy (anonymous):

\[\sqrt[3]{y ^{10}} \sqrt[3]{16y ^{11}}\]

OpenStudy (anonymous):

You can combine the two radicals and simplify...\[\sqrt[3]{16*y^{21}}=(16*y^{21})^{1/3}=\sqrt[3]{16}*y^7\]

OpenStudy (anonymous):

this one says although the answer is equal to the correct answer it is not in the correct form

OpenStudy (anonymous):

\[2y ^{7} \sqrt[2]{2?}\]

OpenStudy (anonymous):

nzxt i posted the answer to the other one for you

OpenStudy (anonymous):

dj.atomix has the answer but missed a part \[\sqrt[3]{16}=\sqrt[3]{8\times 2}=2\sqrt[3]2\]

OpenStudy (anonymous):

so your answer is \[2y^7\sqrt[3]{2}\]

OpenStudy (anonymous):

I was never good at finding those quickly, thanks satellite and mrjoselson for spotting that.

OpenStudy (anonymous):

mrjoselson was correct but for some reason he turned the cube root into the square root and that is an error